# Breaking the Visual Shortcuts in Multimodal Knowledge-Based Visual Question Answering

### v1

@@ Abstract
Existing Multimodal Knowledge-Based Visual Question Answering (MKB-VQA) benchmarks suffer from “visual shortcuts”, as the query image typically matches the primary subject entity of the target document. We demonstrate that models can exploit these shortcuts, achieving comparable results using visual cues alone. To address this, we introduce Relational Entity Text-Image kNowledge Augmented (RETINA) benchmark, automatically constructed using an LLM-driven pipeline, consisting of 120k training and 2k human-curated test set. RETINA contains queries referencing secondary subjects (i.e. related entities) and pairs them with images of these related entities, removing the visual shortcut. When evaluated on RETINA existing models show significantly degraded performance, confirming their reliance on the shortcut. Furthermore, we propose Multi-Image MultImodal Retriever (MIMIR), which enriches document embeddings by augmenting images of multiple related entities, effectively handling RETINA, unlike prior work that uses only a single image per document. Our experiments validate the limitations of existing benchmarks and demonstrate the effectiveness of RETINA and MIMIR. Our project is available at: Project Page.
@@ 1 Introduction
Knowledge-Based Visual Question Answering (KB-VQA) is a challenging task where not only the visual input and its corresponding natural language query are processed, but external world knowledge must be integrated and reasoned over to correctly determine the answer. While early KB-VQA approaches were reliant on unimodal, text-only KBs [chen2023can, mensink2023encyclopedic, marino2019okvqavisualquestionanswering, schwenk2022aokvqabenchmarkvisualquestion, lin2024preflmr, lin2023finegrainedlateinteractionmultimodalretrieval], recent research [deng-etal-2025-muka, caffagni2025recurrence, wei2024uniir, lin2025mmembeduniversalmultimodalretrieval, yu2025cafeunifyingrepresentationgeneration, liu2024lamralargemultimodalmodel, kim2025genius] is shifting to the development of Multimodal KB-VQA (MKB-VQA) to integrate richer information.
@@
In the real world, information relevant to a query is frequently found not in documents directly about the query image but in other related documents. For instance, when asking “Which beetle, native to North America, feeds on this particular plant?” with a photo of an potato, the answer “Lema daturaphila” may be found not in the “Potato” document but rather in the “Lema daturaphila” document, as shown in Fig. 1b. As in such cases, it is often uncertain which document contains the correct answer. However, existing MKB-VQA benchmarks are typically limited to scenarios where each query’s image entity directly matches the primary subject (i.e., main entity) of the target document and its image, leading to a visual shortcut. As illustrated in  Fig. 1a, given a query “In which part of the world does this animal live?” accompanied by an image of the beetle “Lema daturaphila”, existing benchmarks allow trivial retrieval by exploiting visual cues that directly match the main entity. Consequently, models trained on such data tend to rely on these superficial visual correlations. To verify this shortcut, a preliminary experiment where a model [deng-etal-2025-muka] was trained without using the textual component of the query was conducted. As shown in Fig. 2, the model achieved comparable retrieval performance using only the query image, indicating that existing MKB-VQA benchmarks are largely solvable through visual shortcut exploitation.
@@
To address this, we construct Relational Entity Text-Image kNowledge Augmented (RETINA) bench, a MKB-VQA benchmark designed to eliminate such shortcuts and better reflect real-world. RETINA is built through an automated, LLM-driven pipeline that (1) identifies entities related to each document’s main entity (i.e. related entity), (2) constructs a structured knowledge graph, (3) extracts subgraphs for query generation, and (4) generates queries guided by those subgraphs. As prompt-based approaches that feed full documents to LLMs often produce hallucinated or incorrect entity references, we instead employ a graph-guided generation process [ainslie2023gqa, tran2025reasonvqa, trivedi2022musiquemultihopquestionssinglehop, reddy-etal-2017-generating]. This strategy allows controlled query construction by extracting relevant subgraphs corresponding to specific related entities. The resulting queries are then paired with images of the corresponding related, but non-main entities, removing the visual shortcut. Our automated pipeline produces 120K generated training samples alongside a 2K human-verified test set. When evaluating existing models  [lin2024preflmr, deng-etal-2025-muka, wei2024uniir, caffagni2025recurrence], which were trained on shortcut-prone benchmarks, we observe a substantial performance drop on RETINA, reflecting the models’ design that typically augments each document with a single image of its main entity.
@@
To overcome this limitation, we propose Multi-Image MultImodal Retriever (MIMIR) that enriches document representations by integrating images of multiple related entities, extending the single-image framework of prior work [deng-etal-2025-muka]. Specifically, MIMIR extracts related entities from document text, augments multiple images of the related entities, which guides the retrieval of the corresponding documents that contain the related entities, thereby enabling retrieval in the absence of visual shortcuts. We summarize our contributions as follows:
@@
Figure 2: Preliminary Experiment on Existing Benchmarks. Recall@$k$ on benchmarks such as Infoseek [chen2023can] and EVQA [mensink2023encyclopedic], which exhibit visual shortcuts [deng-etal-2025-muka]. Models fine-tuned with image-only (blue) queries versus image+text (green) queries.
@@ 2 Related Works
Knowledge-based VQA Benchmarks Early developments in KB-VQA [lerner2022viquae, lin2024preflmr, hu2023avis, lin2023fine, marino2019okvqavisualquestionanswering, schwenk2022aokvqabenchmarkvisualquestion, jain2021select, wang2017fvqa, shahMYP19] incorporated external textual KBs such as Wikipedia, enabling responses grounded in external knowledge. Later developments [deng-etal-2025-muka, caffagni2025recurrence, wei2024uniir] expanded the textual KBs into the multimodal KBs by associating each textual document with corresponding images. However, these existing MKB-VQA benchmarks are biased towards samples with visual shortcuts, where the query’s image depicts the same visual entity as the document’s image entity. To address this bias, we introduce RETINA, a new benchmark that explicitly eliminates such shortcuts.
Multimodal Retrievers Early multimodal retrievers  [radford2021learning, Mori1999ImagetowordTB, kolesnikov2020bigtransferbitgeneral] primarily focused on image-to-text and text-to-image retrieval by aligning visual and textual representations within a shared embedding space. Subsequent studies  [Gao_2022_CVPR, lin2023finegrainedlateinteractionmultimodalretrieval, lin2024preflmr, long2025retrieval, lin2022revive] extended this paradigm to handle queries that jointly incorporate both textual and visual inputs. Recent multimodal KB retrieval approaches [deng-etal-2025-muka, caffagni2025recurrence] represent documents with both textual and visual modalities and demonstrate strong performance on tasks involving fine-grained, knowledge-intensive entities. In parallel, universal retrieval frameworks have been proposed [wei2024uniir, yu2025cafeunifyingrepresentationgeneration, lin2025mmembeduniversalmultimodalretrieval, kim2025genius, liu2025lamra] enabling retrieval across arbitrary query and document modality pairs. These existing retrievers encode each document with a single image, but real-world queries often involve diverse entities. Our MIMIR employs multi-image encoding to better handle realistic multimodal queries.
@@ 3 Dataset Construction
Figure 3: RETINA Benchmark Generation Pipeline. (a) constructs one-hop neighborhood graphs by extracting named entities (gray box) related to the answer entity (white box) and their relations using an LLM [bai2023qwen]; (b) samples an query entity (green) and qualifying entity (teal) to form a target subgraph with the answer entity (red); and (c) feeds the target subgraph into an LLM to generate a textual query and collect a corresponding image from M2KR Images [lin2024preflmr]. The query is then paraphrased to minimize lexical overlap with the document.
Our goal is to build an MKB-VQA benchmark that captures challenging real-world scenarios, specifically those where visual shortcuts are absent. In practice, information associated with a particular entity is often scattered across multiple documents rather than confined to the one describing that entity itself. Consequently, user are inherently uncertain about which document contains the answer to a given query. However, existing MKB-VQA benchmarks are biased toward scenarios where the main entity of the target document typically has the same entity as the query image. To address this bias, we focus on collecting samples that do not permit such visual shortcuts, by deliberately selecting a query image that differ from the main entity of the target Wikipedia document. Specifically, we employ an LLM [bai2023qwen] to identify entities related to the main entity within the target document, and generate a multimodal query by selecting the query image from these related entities. Through this automatic generation pipeline, we construct a large-scale benchmark RETINA, consisting of 120K training samples and a 2K human-curated test set. When combined with existing benchmarks, our RETINA bench provides a realistic evaluation setting for MKB-VQA in real-world scenarios.
@@ 3.1 MKB-VQA Sample Generation Pipeline
Automatically generating VQA samples from raw Wikipedia documents is challenging for LLMs because the documents are often long and the relationships between entities are implicit and scattered throughout the text [liu2023lostmiddlelanguagemodels]. We address this by explicitly extracting relationships between entities from the documents before generating VQA samples. Specifically, we adopt a three-stage pipeline: (a) converting the document into a knowledge graph to enable better control over the generation process, (b) selecting target subgraphs, and (c) generating multimodal QA based on the selected subgraphs.
@@
LLM-based VQA Sample Generation Given an extracted target subgraph, our goal is to generate a QA pair for MKB-VQA. We first obtain a query image by retrieving an image of the query entity from M2KR [lin2024preflmr]. Then, we prompt the LLM to formulate a question whose answer should correspond to the answer entity, conditioned on two natural language descriptions describing the relations among the entities in the subgraph. To prevent textual leakage, we explicitly instruct the LLM not to mention the query entity by name within the question, but instead to refer to it using an appropriate demonstrative phrase (e.g., “this organization” or “this monument”), ensuring that the model must infer it from the query image rather than the text. Finally, the generated question is passed to the LLM once more for paraphrasing, and we filter out questions that can be retrieved by BM25 [robertson1995okapi] in the top-5 results, thereby reducing lexical overlap with the source document. The full prompts are provided in the Supp. Mat., and Fig. 3c) illustrates this process along with an example QA pair.
@@ 3.2 RETINA Bench Construction
Figure 4: Overview of MIMIR Document Encoder Architecture. (a) given a document, related named entities are identified with an LLM [bai2023qwen], and corresponding images are collected from the KB; (b) textual, global image, and patch-level features are extracted, with patch features attending to textual features through cross-attention to yield multimodal features; and (c) entity token embeddings are incorporated into the textual features prior to cross-attention for richer contextualization; (d) the final document embedding jointly integrates textual, global, and multimodal features projected into the same embedding space.
@@ 4 Method
Existing multimodal retrievers [deng-etal-2025-muka, caffagni2025recurrence, wei2024uniir] typically encode each document together with an image depicting its main entity. While these approaches achieve strong performance on existing MKB-VQA benchmarks, their effectiveness diminishes on RETINA. This limitation arises because each document embedding is augmented solely with the image of its main entity. Consequently, these systems often fail to retrieve the correct document when the query image depicts a non-main entity within the target document. To address this issue, we introduce MIMIR, a multimodal retriever that enriches document representations by incorporating images of multiple related entities. Our approach builds upon MuKA [deng-etal-2025-muka], a multimodal single-image document retriever, and extends its representation capacity. In the following, we first review the scoring mechanism and query encoder of [deng-etal-2025-muka] (Sec. 4.1 and 4.2), and then describe our extended document embedding module in detail (Sec. 4.3).
@@ 4.1 Late Interaction Scoring
Following ColBERT family of multimodal retrievers [deng-etal-2025-muka, lin2024preflmr], we represent both the query and document as sets of token-level vector, $\mathcal{Q}=\{q_{1},q_{2},\dots,q_{m}\}$ and $\mathcal{D}=\{d_{1},d_{2},\dots,d_{n}\}$, respectively. Each query token vector $q_{i}$ interacts with every document token vector in $\mathcal{D}$, and the contribution of the query token to the overall relevance score is defined by the maximum similarity across document tokens. The overall relevance score is then computed by summing these maximum contributions across all query tokens. Formally, the document’s overall score is computed as
@@
where $\text{sim}(q_{i},d_{j})$ denotes the cosine similarity between query token $q_{i}$ and document token $d_{j}$. Despite involving complex token-wise similarity computations, practical efficiency is supported by an efficient indexing following [santhanam2022colbertv2effectiveefficientretrieval].
@@ 4.2 Query Encoder
The query encoder comprises image and text encoders and a transformer block with a cross-attention layer, each paired with its own projection layers. Given a user query composed of an image and text $(I^{(q)},T^{(q)})$, the goal is to produce the query feature set $\mathcal{Q}$ for subsequent scoring. The textual and visual inputs are first processed separately by a ColBERTv2 text encoder [santhanam2022plaidefficientenginelate] and a CLIP image encoder [radford2021learning], yielding a set of $N_{t}$ token-level text embeddings $\mathcal{T}^{(q)}=\{t^{(q)}_{1},t^{(q)}_{2},\dots,t^{(q)}_{N_{t}}\}$ and a global image feature $g^{(q)}$, respectively. In addition, a set of patch-level features $\mathcal{P}^{(q)}=\{p^{(q)}_{1},p^{(q)}_{2},\dots,p^{(q)}_{N_{p}}\}$ is extracted from the penultimate layer of the CLIP encoder, where $N_{p}=H\times W$ denotes the number of spatial patches. These patch features are then contextualized using the textual features $\mathcal{T}^{(q)}$ through the transformer block with cross-attention [vaswani2023attentionneed]. Specifically, $\mathcal{P}^{(q)}$ serves as the queries, while $\mathcal{T}^{(q)}$ provides the keys and values, producing a set of $N_{p}$ multimodal features $\mathcal{M}^{(q)}=\{m^{(q)}_{1},m^{(q)}_{2},\dots,m^{(q)}_{N_{p}}\}$. Finally, the query feature set $\mathcal{Q}$ is formed by projecting the global image feature $g^{(q)}$, all textual features $t^{(q)}_{i}$, and all multimodal features $m^{(q)}_{j}$ into $\mathbb{R}^{128}$ using their dedicated projection layers. Denoting the projected features by $\widehat{g}^{(q)}$, $\widehat{\mathcal{T}}^{(q)}=\{\widehat{t}^{(q)}_{i}\}$, and $\widehat{\mathcal{M}}^{(q)}=\{\widehat{m}^{(q)}_{j}\}$, respectively, the final query feature set is $\mathcal{Q}=\{\widehat{g}^{(q)}\}\cup\widehat{\mathcal{T}}^{(q)}\cup\widehat{\mathcal{M}}^{(q)}$ where all elements lie in $\mathbb{R}^{128}$.

### v2

@@ Abstract
Existing Multimodal Knowledge-Based Visual Question Answering (MKB-VQA) benchmarks suffer from “visual shortcuts”, as the query image typically matches the primary subject entity of the target document. We demonstrate that models can exploit these shortcuts, achieving comparable results using visual cues alone. To address this, we introduce Relational Entity Text-Image kNowledge Augmented (RETINA) benchmark, automatically constructed using an LLM-driven pipeline, consisting of 120k training and 2k human-curated test set. RETINA contains queries referencing secondary subjects (i.e. related entities) and pairs them with images of these related entities, removing the visual shortcut. When evaluated on RETINA existing models show significantly degraded performance, confirming their reliance on the shortcut. Furthermore, we propose Multi-Image MultImodal Retriever (MIMIR), which enriches document embeddings by augmenting images of multiple related entities, effectively handling RETINA, unlike prior work that uses only a single image per document. Our experiments validate the limitations of existing benchmarks and demonstrate the effectiveness of RETINA and MIMIR.
@@ 1 Introduction
Knowledge-Based Visual Question Answering (KB-VQA) is a challenging task where not only the visual input and its corresponding natural language query are processed, but external world knowledge must be integrated and reasoned over to correctly determine the answer. While early KB-VQA approaches were reliant on unimodal, text-only KBs [5, 27, 26, 35, 19, 18], recent research [6, 4, 40, 16, 42, 23, 12] is shifting to the development of Multimodal KB-VQA (MKB-VQA) to integrate richer information.
@@
In the real world, information relevant to a query is frequently found not in documents directly about the query image but in other related documents. For instance, when asking “Which beetle, native to North America, feeds on this particular plant?” with a photo of an potato, the answer “Lema daturaphila” may be found not in the “Potato” document but rather in the “Lema daturaphila” document, as shown in Fig. 1b. As in such cases, it is often uncertain which document contains the correct answer. However, existing MKB-VQA benchmarks are typically limited to scenarios where each query’s image entity directly matches the primary subject (i.e., main entity) of the target document and its image, leading to a visual shortcut. As illustrated in  Fig. 1a, given a query “In which part of the world does this animal live?” accompanied by an image of the beetle “Lema daturaphila”, existing benchmarks allow trivial retrieval by exploiting visual cues that directly match the main entity. Consequently, models trained on such data tend to rely on these superficial visual correlations. To verify this shortcut, a preliminary experiment where a model [6] was trained without using the textual component of the query was conducted. As shown in Fig. 2, the model achieved comparable retrieval performance using only the query image, indicating that existing MKB-VQA benchmarks are largely solvable through visual shortcut exploitation.
@@
To address this, we construct Relational Entity Text-Image kNowledge Augmented (RETINA) bench, a MKB-VQA benchmark designed to eliminate such shortcuts and better reflect real-world. RETINA is built through an automated, LLM-driven pipeline that (1) identifies entities related to each document’s main entity (i.e. related entity), (2) constructs a structured knowledge graph, (3) extracts subgraphs for query generation, and (4) generates queries guided by those subgraphs. As prompt-based approaches that feed full documents to LLMs often produce hallucinated or incorrect entity references, we instead employ a graph-guided generation process [1, 36, 37, 30]. This strategy allows controlled query construction by extracting relevant subgraphs corresponding to specific related entities. The resulting queries are then paired with images of the corresponding related, but non-main entities, removing the visual shortcut. Our automated pipeline produces 120K generated training samples alongside a 2K human-verified test set. When evaluating existing models  [19, 6, 40, 4], which were trained on shortcut-prone benchmarks, we observe a substantial performance drop on RETINA, reflecting the models’ design that typically augments each document with a single image of its main entity.
@@
To overcome this limitation, we propose Multi-Image MultImodal Retriever (MIMIR) that enriches document representations by integrating images of multiple related entities, extending the single-image framework of prior work [6]. Specifically, MIMIR extracts related entities from document text, augments multiple images of the related entities, which guides the retrieval of the corresponding documents that contain the related entities, thereby enabling retrieval in the absence of visual shortcuts. We summarize our contributions as follows:
@@
Figure 2: Preliminary Experiment on Existing Benchmarks. Recall@$k$ on benchmarks such as Infoseek [5] and EVQA [27], which exhibit visual shortcuts [6]. Models fine-tuned with image-only (blue) queries versus image+text (green) queries.
@@ 2 Related Works
Knowledge-based VQA Benchmarks Early developments in KB-VQA [14, 19, 10, 17, 26, 35, 11, 39, 32] incorporated external textual KBs such as Wikipedia, enabling responses grounded in external knowledge. Later developments [6, 4, 40] expanded the textual KBs into the multimodal KBs by associating each textual document with corresponding images. However, these existing MKB-VQA benchmarks are biased towards samples with visual shortcuts, where the query’s image depicts the same visual entity as the document’s image entity. To address this bias, we introduce RETINA, a new benchmark that explicitly eliminates such shortcuts.
Multimodal Retrievers Early multimodal retrievers  [29, 28, 13] primarily focused on image-to-text and text-to-image retrieval by aligning visual and textual representations within a shared embedding space. Subsequent studies  [7, 18, 19, 25, 20] extended this paradigm to handle queries that jointly incorporate both textual and visual inputs. Recent multimodal KB retrieval approaches [6, 4] represent documents with both textual and visual modalities and demonstrate strong performance on tasks involving fine-grained, knowledge-intensive entities. In parallel, universal retrieval frameworks have been proposed [40, 42, 16, 12, 24] enabling retrieval across arbitrary query and document modality pairs. These existing retrievers encode each document with a single image, but real-world queries often involve diverse entities. Our MIMIR employs multi-image encoding to better handle realistic multimodal queries.
@@ 3 Dataset Construction
Figure 3: RETINA Benchmark Generation Pipeline. (a) constructs one-hop neighborhood graphs by extracting named entities (gray box) related to the answer entity (white box) and their relations using an LLM [2]; (b) samples an query entity (green) and qualifying entity (teal) to form a target subgraph with the answer entity (red); and (c) feeds the target subgraph into an LLM to generate a textual query and collect a corresponding image from M2KR Images [19]. The query is then paraphrased to minimize lexical overlap with the document.
Our goal is to build an MKB-VQA benchmark that captures challenging real-world scenarios, specifically those where visual shortcuts are absent. In practice, information associated with a particular entity is often scattered across multiple documents rather than confined to the one describing that entity itself. Consequently, user are inherently uncertain about which document contains the answer to a given query. However, existing MKB-VQA benchmarks are biased toward scenarios where the main entity of the target document typically has the same entity as the query image. To address this bias, we focus on collecting samples that do not permit such visual shortcuts, by deliberately selecting a query image that differ from the main entity of the target Wikipedia document. Specifically, we employ an LLM [2] to identify entities related to the main entity within the target document, and generate a multimodal query by selecting the query image from these related entities. Through this automatic generation pipeline, we construct a large-scale benchmark RETINA, consisting of 120K training samples and a 2K human-curated test set. When combined with existing benchmarks, our RETINA bench provides a realistic evaluation setting for MKB-VQA in real-world scenarios.
@@ 3.1 MKB-VQA Sample Generation Pipeline
Automatically generating VQA samples from raw Wikipedia documents is challenging for LLMs because the documents are often long and the relationships between entities are implicit and scattered throughout the text [22]. We address this by explicitly extracting relationships between entities from the documents before generating VQA samples. Specifically, we adopt a three-stage pipeline: (a) converting the document into a knowledge graph to enable better control over the generation process, (b) selecting target subgraphs, and (c) generating multimodal QA based on the selected subgraphs.
@@
LLM-based VQA Sample Generation Given an extracted target subgraph, our goal is to generate a QA pair for MKB-VQA. We first obtain a query image by retrieving an image of the query entity from M2KR [19]. Then, we prompt the LLM to formulate a question whose answer should correspond to the answer entity, conditioned on two natural language descriptions describing the relations among the entities in the subgraph. To prevent textual leakage, we explicitly instruct the LLM not to mention the query entity by name within the question, but instead to refer to it using an appropriate demonstrative phrase (e.g., “this organization” or “this monument”), ensuring that the model must infer it from the query image rather than the text. Finally, the generated question is passed to the LLM once more for paraphrasing, and we filter out questions that can be retrieved by BM25 [31] in the top-5 results, thereby reducing lexical overlap with the source document. The full prompts are provided in the Supp. Mat., and Fig. 3c) illustrates this process along with an example QA pair.
@@ 3.2 RETINA Bench Construction
Figure 4: Overview of MIMIR Document Encoder Architecture. (a) given a document, related named entities are identified with an LLM [2], and corresponding images are collected from the KB; (b) textual, global image, and patch-level features are extracted, with patch features attending to textual features through cross-attention to yield multimodal features; and (c) entity token embeddings are incorporated into the textual features prior to cross-attention for richer contextualization; (d) the final document embedding jointly integrates textual, global, and multimodal features projected into the same embedding space.
@@ 4 Method
Existing multimodal retrievers [6, 4, 40] typically encode each document together with an image depicting its main entity. While these approaches achieve strong performance on existing MKB-VQA benchmarks, their effectiveness diminishes on RETINA. This limitation arises because each document embedding is augmented solely with the image of its main entity. Consequently, these systems often fail to retrieve the correct document when the query image depicts a non-main entity within the target document. To address this issue, we introduce MIMIR, a multimodal retriever that enriches document representations by incorporating images of multiple related entities. Our approach builds upon MuKA [6], a multimodal single-image document retriever, and extends its representation capacity. In the following, we first review the scoring mechanism and query encoder of [6] (Sec. 4.1 and 4.2), and then describe our extended document embedding module in detail (Sec. 4.3).
@@ 4.1 Late Interaction Scoring
Following ColBERT family of multimodal retrievers [6, 19], we represent both the query and document as sets of token-level vector, $\mathcal{Q}=\{q_{1},q_{2},\dots,q_{m}\}$ and $\mathcal{D}=\{d_{1},d_{2},\dots,d_{n}\}$, respectively. Each query token vector $q_{i}$ interacts with every document token vector in $\mathcal{D}$, and the contribution of the query token to the overall relevance score is defined by the maximum similarity across document tokens. The overall relevance score is then computed by summing these maximum contributions across all query tokens. Formally, the document’s overall score is computed as
@@
where $\text{sim}(q_{i},d_{j})$ denotes the cosine similarity between query token $q_{i}$ and document token $d_{j}$. Despite involving complex token-wise similarity computations, practical efficiency is supported by an efficient indexing following [34].
@@ 4.2 Query Encoder
The query encoder comprises image and text encoders and a transformer block with a cross-attention layer, each paired with its own projection layers. Given a user query composed of an image and text $(I^{(q)},T^{(q)})$, the goal is to produce the query feature set $\mathcal{Q}$ for subsequent scoring. The textual and visual inputs are first processed separately by a ColBERTv2 text encoder [33] and a CLIP image encoder [29], yielding a set of $N_{t}$ token-level text embeddings $\mathcal{T}^{(q)}=\{t^{(q)}_{1},t^{(q)}_{2},\dots,t^{(q)}_{N_{t}}\}$ and a global image feature $g^{(q)}$, respectively. In addition, a set of patch-level features $\mathcal{P}^{(q)}=\{p^{(q)}_{1},p^{(q)}_{2},\dots,p^{(q)}_{N_{p}}\}$ is extracted from the penultimate layer of the CLIP encoder, where $N_{p}=H\times W$ denotes the number of spatial patches. These patch features are then contextualized using the textual features $\mathcal{T}^{(q)}$ through the transformer block with cross-attention [38]. Specifically, $\mathcal{P}^{(q)}$ serves as the queries, while $\mathcal{T}^{(q)}$ provides the keys and values, producing a set of $N_{p}$ multimodal features $\mathcal{M}^{(q)}=\{m^{(q)}_{1},m^{(q)}_{2},\dots,m^{(q)}_{N_{p}}\}$. Finally, the query feature set $\mathcal{Q}$ is formed by projecting the global image feature $g^{(q)}$, all textual features $t^{(q)}_{i}$, and all multimodal features $m^{(q)}_{j}$ into $\mathbb{R}^{128}$ using their dedicated projection layers. Denoting the projected features by $\widehat{g}^{(q)}$, $\widehat{\mathcal{T}}^{(q)}=\{\widehat{t}^{(q)}_{i}\}$, and $\widehat{\mathcal{M}}^{(q)}=\{\widehat{m}^{(q)}_{j}\}$, respectively, the final query feature set is $\mathcal{Q}=\{\widehat{g}^{(q)}\}\cup\widehat{\mathcal{T}}^{(q)}\cup\widehat{\mathcal{M}}^{(q)}$ where all elements lie in $\mathbb{R}^{128}$.